OpenStudy (anonymous):

find the discriminant and simplified radical roots of the following: 2x^2 + 11x + 5

OpenStudy (anonymous):

Do you know discriminant formula?

OpenStudy (anonymous):

\[ax^2 + bx + c\] discriminant: \[b^2 - 4ac\] finding roots there are a few different ways but since you have already found the discriminant: \[x = \frac{-b \pm \sqrt{b^2-4ac}}{2a}\]

OpenStudy (anonymous):

if discriminant < 0 then there are no real roots , if discriminant is a perfect square then the equation factorises into (Ax+B)(Cx+D) form

OpenStudy (anonymous):

so do i just keep it in the form of (-11 ±√81)/4

Directrix (directrix):

Discriminant: b^2 -4ac = 11^2 - 4 (2)(5) = 121 - 40 = 81. x = (-11 ±√81 )/ 4 = ( -11 + 9) / 4 = -2/4 = -1/2 OR x = ( -11 - 9) / 4 = -20/4 = -5. Roots are: -1/2 and -5

OpenStudy (anonymous):

@directrix okay so thats what i did before. i just thought since it said radical roots that i keep it in radical form. so if i were given the equation x^2 - 63 the roots would be kept in radical for right? x= ±3√7

Directrix (directrix):

Yes, those are in simplified radical form. In the first problem, √ 81 was in radical form but not simplified radical form because √ 81 = 9.

Directrix (directrix):

Minor Point: equation x^2 - 63--> is not an equation but a quadratic expression. x^2 - 63 = 0 is the equation to find the roots (x-intercepts) of the quadratic equation y = x^2 - 63

OpenStudy (anonymous):

so wait is my answer correct then ? x^2 - 63 = ±3√7 cause i'm asked to find the root of that as well

Directrix (directrix):

What is the equation you were given? I don't see it.

OpenStudy (anonymous):

oh cause in the question it asks me to find the discriminant of 2x^2 + 11x +5 x^2-63, and 5x^2-6x-2

Directrix (directrix):

x^2 - 63 = x^2 + 0x - 63 = 0. discriminant: 0^2 - 4(1)(-63) = 0 + 252 = 252 What did you write for the discriminant of this?

OpenStudy (anonymous):

yeah i got the same discriminant too

Directrix (directrix):

So, were you asked to find the roots also?

OpenStudy (anonymous):

yep. so what i did was after i put into the quadratic formula i got (-0 ± √252)/2 = (± √36 x 7)/2 =(±6√7)/2 =±3√7

OpenStudy (anonymous):

is that right?

Directrix (directrix):

Yes, but you lost the variable: x = ±3√7 This statement is incorrect x^2 - 63 = ±3√7 The roots of x^2 - 63 = 0 are ±3√7 is correct.

OpenStudy (anonymous):

so it must be written as: The roots of x^2 - 63 = 0 are ±3√7

Directrix (directrix):

Yes. Did you get the discriminant and roots of 5x^2-6x-2 ?

OpenStudy (anonymous):

yes i did. the discriminant is 76 and the roots of 5x^2-6x-2 =0 are - 6 ±2√9

Directrix (directrix):

@alexeis_nicole --> I got x = ( 3 ± √19 ) / 5 for the roots of 5x^2-6x-2 =0. You check your work; I'll check mine.

OpenStudy (anonymous):

okay this is what i got. the discriminant is 76 the roots of 5x^2 - 6x -2 x= (6± √76)10 = (6± √4 x 19)10 = (6± 2√19)10 .. ? where did i go wrong ?

Directrix (directrix):

(6± 2√19) / 10 = 2 (3 ± √19 ) /10 = (3 ± √19 ) / 5 --> You didn't simplify the fractions.

OpenStudy (anonymous):

OOOH ! okay. i forgot to simplify. thanks directix :)

Directrix (directrix):

You are welcome.